OpenStudy (anonymous):

Explain, in complete sentences, how you would completely factor 2x^3 - 24x^2y + 12xy - 144y^2 and check your factors for accuracy.

OpenStudy (anonymous):

try pairing the terms: no first divide thru by 2: = 2( x^3 -12x^2y + 6xy - 72y^2 ) now pair terms: = 2(x^3 + 6xy - 12x^2y - 72y^2) factor out = 2[ x(6y +x^2) -12y (6y + x^2)] two brackets are equal so = 2(x - 12y)(6y + x^2)

OpenStudy (anonymous):

to check we expand in order to hopefully get back to origainal expression 2( 6xy + x^3 -72y^2 - 12x^2_ = 2x^3 -24x^2y + 12xy - 144y^2